OpenStudy (anonymous):

Atomic sodium is studied by photoelectron spectroscopy. What is the maximum wavelength of the incident radiation that will ionize the most weakly held electrons in sodium and scatter them so that their de Broglie wavelength measures 3.091e-10 m?

OpenStudy (anonymous):

Try using E=(mv^2)/2 to get the E, and then use E=HC/lambda :)

OpenStudy (anonymous):

I think that's what I did...

OpenStudy (anonymous):

Will you share your answer, please? Time is short :)

OpenStudy (anonymous):

Jajajaja sure, but try to do the calculations on your own so that you can understand it :) 1.12*10^-8

OpenStudy (anonymous):

I did. Head's exploding. And by the way: It's wrong! :)

OpenStudy (anonymous):

really? must have been a different one, cause I gave yu the answer I had for number 9...

OpenStudy (anonymous):

you*... Sorryy

OpenStudy (anonymous):

doesn't matter. I'm over the magic 60% already! :)

OpenStudy (aaronq):

deBroglie wavelength is described by: \[\lambda =\frac{ h }{ mv }\]